A boy runs along a straight path for the first half distance with a velocity u and the second half distance with velocity v . Find the average 

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Let the total distance bes
So the time taken to cover first half =t_1=\dfrac{s/2}{u}
Time taken to cover second half=t_2=\dfrac{s/2}{v}

Hence total time takent=t=t_1+t_2=(\dfrac{s}{2})(\dfrac{1}{u}+\dfrac{1}{v})
Average velocity=\dfrac{s}{t}=\dfrac{2}{\dfrac{1}{u}+\dfrac{1}{v}}